1. Expand the following using the Binomial Theorem



a. (x-3y)^3



b. (2a+b^-3)^4



c. (a+b+c)²



2. Solve the recurrence relation Pn = Pn-1 + n with the initial condition p_1 = 2 using iteration.

1
Expert's answer
2021-11-22T20:17:55-0500

a.


"(x-3y)^3=x^3-9x^2y+27xy^2-27y^3"

b.


"(2a+b^{-3})^4=16a^4+32a^3b^{-3}+24a^2b^{-6}+8ab^{-9}+b^{-12}"

c.


"(a+b+c)^2=a^2+b^2+c^2+2ab+2ac+2bc"

d.


"p_2=2+2=4"

"p_3=4+3=7"

"p_4=7+4=11"

"p_5=11+5=16"

"\vdots"

"(p_2-p_1)+(p_3-p_2)+(p_4-p_3)+..."

"+(p_{n-1}-p_{n-2})+(p_n-p_{n-1})=2+3+4+..."

"+(n-1)+n"

"p_n-p_1=\displaystyle\sum_{i=1}^ni-1"

"p_n-p_1=\dfrac{n(n+1)}{2}-1"

"p_n=1+\dfrac{n(n+1)}{2}"